How far (in cm) would light travel through diamond in 1.00 μs?




1
Expert's answer
2021-11-14T17:15:39-0500

The speed of light in diamond is approximately


"v= \dfrac{c}{n_d} = \dfrac{3\times 10^{10}cm/s}{2.42} \approx 1.24\times 10^{10}cm/s"


(here "n_d = 2.42" is the refractive index of diamond).

Thus, in time "t = 1\mu s = 10^{-6}s" it will travel:


"d = vt = 1.24\times 10^{10}cm/s\cdot 10^{-6}s=1.24\times 10^{4}cm"

Answer. "1.24\times 10^{4}cm".